What is the solubility of La(IO₃)₃ in a solution that contains 0.300 M La³⁺ ions? (Ksp of La(IO₃)₃ is 7.5 × 10⁻¹²) Please show all steps.
Answer #1
  1. La (103) - Låt + 3 loj 0.3 +S 0.3+) +35 35 0.3+5= 0.3 ksp= ( La 3r] [ Ig]3 7.5x102= (0.3+3) (35) 2 953 = 2.5x10-1 s32 9.26 X